If the hypotenuses a square is 6 Cm, what is the length of the side? Show all work
Answer by rothauserc(4718)

each side of a square is of equal length, therefore
s^2 + s^2 = 36
2s^2 = 36
s^2 = 18
s = square root(18) = 3*square root(2) = 4.242640687 cm